51. Suppose that a firm has the production function

Q = 2K1/2 L1/2

Assume that the capital stock is fixed at 9 units (i.e. K =9). If the price of output (P) is $6 per unit, and the wage rate (w) is $2 per unit, determine the optimal or profit-maximizing rate of labor to be hired. What labor-input rate is optimal if the wage rate increased to $3 per unit?
1
Expert's answer
2020-11-16T07:29:01-0500

"Q=2*9^{1\/2}*L^{1\/2}"

"L=(\frac{Q}{6})^2"

"C=9+L*2" ; C-cost

"C=9+\frac{Q^2}{18}" ; "E=PQ=6Q"; E-earnings

"I=E-C"; I-Income

the quantity that gives max income  :

"\frac{dI}{dQ}=(6Q-9-\frac{Q^2}{18})'=6-\frac{2*Q*18}{18^2}=6-\frac{2Q}{18}"

"0=6-\frac{2Q}{18}" ; "Q=54"; "L=81"

W=3:

"C=9+3*L"; "C=9+\frac{Q^2}{12}"

"0=6-\frac{2Q}{12}"; "Q=36"

"L=36"